OpenStudy (anonymous):

The table shows the results of drawing colored marbles from a bag. Based on theoretical probability, which is the best prediction for the number of times a red marble would be drawn in 1000 trials? A. 100 B. 125 C. 167 D. 200

OpenStudy (anonymous):

here is the chart

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

cool pic

OpenStudy (anonymous):

cool pic @redbullx and @noseboy908

OpenStudy (noseboy908):

I'll help, I'm just looking at the chart now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so can u figure it out

OpenStudy (noseboy908):

Ok, firstly, add up the total number of marbles present in your test. Do this by adding up the FIRST column. Tell me what you get.

OpenStudy (anonymous):

40

OpenStudy (anonymous):

right

OpenStudy (noseboy908):

No, try again.

OpenStudy (anonymous):

do i add the actucul # of marbels and # of marbels drawn

OpenStudy (noseboy908):

No, ONLY the FIRST column.

OpenStudy (anonymous):

ok so add 5+10=15 right

OpenStudy (noseboy908):

yes, as well as ALL the numbers in that column alone.

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

And what is the total?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

1000 times the fraction of red marbles

OpenStudy (anonymous):

i got 100 @noseboy908

OpenStudy (anonymous):

the fraction of red marbles is 5/40 = 1/8

OpenStudy (noseboy908):

Add up the numbers from the FIRST column.

OpenStudy (anonymous):

it is 15

OpenStudy (noseboy908):

You see the column labeled "actual number of marbles"? Add up all the numbers below it.

OpenStudy (anonymous):

what is the relative frequency of the red marbles? did you learn that term?

OpenStudy (anonymous):

and how to compute it?

OpenStudy (anonymous):

i got 40 @noseboy908

OpenStudy (anonymous):

so there are a total of 40 marbles. how many of them are red?

OpenStudy (noseboy908):

Yep, @pgpilot326 is correct.

OpenStudy (anonymous):

15 @pgpilot326

OpenStudy (anonymous):

15? where did you get that?

OpenStudy (noseboy908):

According to the "actual number of marbles" column, what is the number next to "Red"?

OpenStudy (anonymous):

i mean 10 or 5

OpenStudy (anonymous):

number of actual is 5 and number of pulled out is 10

OpenStudy (anonymous):

@Spunkeymonkey822 just ignore the last column. it has nothing to do with the theoretical values.

OpenStudy (anonymous):

ok i just need answer plz

OpenStudy (anonymous):

my test is almost done

OpenStudy (anonymous):

i provided you with all the info to compute. look at my first and second posts and put it together. make sense of it.

OpenStudy (anonymous):

but what u mean is that i add red in both colums

OpenStudy (anonymous):

no, ignore the last column. of the 40 marbles in the bag, 5 of them are red. thus the fraction of red marbles is \(\frac{5}{40}=\frac{1}{8}\)